OpenStudy (anonymous):

] Solve the following system of equations. 3x - 2y = 6 6x - 4y = 14

OpenStudy (mathteacher1729):

What method do they want you to use: 1. graphing 2. substitution 3. elimination 4. other? What methods have you already tried?

OpenStudy (anonymous):

i did substitution and got no solution

OpenStudy (mathteacher1729):

That makes sense. Here's a shortcut way of looking at it... 3x - 2y = 6 6x - 4y = 14 divide second equation by two on both sides, leave the first alone: 3x - 2y = 6 3x - 2y = 8 Uhohs. This is saying that 6 = 8 (Do you know why?) Clearly 6 does not equal 8, and there are no solutions. :)